# A simple approach to rotationally invariant machine learning of a vector quantity

### v1

@@ I Introduction
There are several approaches in the literature to handle orientation dependence. For dipole moments, neural networks were suggested that learn atomic charges to reconstruct the dipole moments as accurately as possible(Unke and Meuwly, 2019; Chen et al., 2024a). For polarizabilities and nonadiabatic couplings, such neural networks were built that can learn the virtual quantities, the analytical derivatives of which produce the required tensorial propertiesWestermayr et al. (2020b); Zhang et al. (2020a); Chen et al. (2024a). The latter approach is inspired by (MLPs) where it is common to derive energy gradients (negative forces) as the derivatives of the machine learning function of potential energies. Many approaches rely on building equivariance into the design of the ML model, either into the local descriptors employed in neural network (NN) approaches, or in the design of the NN itself. This was done in case of Gaussian process regression (GPR), by starting with a kernel based on a smooth overlap of atomic positions (SOAP)(Bartók et al., 2013) and deriving a hierarchy of -SOAP kernels preserving rotational symmetry(Glielmo et al., 2017; Grisafi et al., 2018; Wilkins et al., 2019). This so-called symmetry-adapted GPR (SA-GPR) approach was used to paracetamol system leading to a covariant predictions(Raimbault et al., 2019). The same is possible for the class of ML algorithms based on NNs as shown for neuroevolution potential (NEP) scheme(Fan et al., 2021), which was modified into tensorial-NEP approach(Xu et al., 2024) or embedded atom NN (EANN)(Zhang et al., 2019) transformed into tensorial EANN,(Zhang et al., 2020a) outperforming SA-GPR. Other examples introduced new rotationally covariant architectures(Anderson et al., 2019), equivariant message passing NN(Schütt et al., 2021) or tensor field NN(Thomas et al., 2018). The NN-based approaches proved to be excellent tools for predicting infrared and Raman spectra(Chen et al., 2024b; Zhang et al., 2020b), even under the influence of external fields(Zhang and Jiang, 2023).
When considering the kernel ridge regression (KRR) together with a global molecular descriptor, an alternative to these previous approaches is to rotate the molecular geometry to some reference frame(Bereau et al., 2015; Liang et al., 2017). Although such a rotation is straightforward to implement, it cannot be used for flexible or dissociative systems(Grisafi et al., 2018). Here we built upon this rotational idea to propose a three-step procedure to maintain proper covariance. We demonstrate this new method (named rotate-predict-rotate for reason that will be clear in the next section) on 1,2-dichloroethane, for which new descriptors were introduced improving the overall test set root mean square error (RMSE) values for both dipole moment and polarizability components.

### v2

@@ I Introduction
There are several approaches in the literature to handle orientation dependence. For dipole moments, neural networks were suggested that learn atomic charges to reconstruct the dipole moments as accurately as possible(Unke and Meuwly, 2019; Chen et al., 2024a). For polarizabilities and nonadiabatic couplings, such neural networks were built that can learn the virtual quantities, the analytical derivatives of which produce the required tensorial propertiesWestermayr et al. (2020b); Zhang et al. (2020a); Chen et al. (2024a). The latter approach is inspired by MLPs where it is common to derive energy gradients (negative forces) as the derivatives of the machine learning function of potential energies. Many approaches rely on building equivariance into the design of the ML model, either into the local descriptors employed in neural network (NN) approaches, or in the design of the NN itself. This was done in case of Gaussian process regression (GPR), by starting with a kernel based on a smooth overlap of atomic positions (SOAP)(Bartók et al., 2013) and deriving a hierarchy of -SOAP kernels preserving rotational symmetry(Glielmo et al., 2017; Grisafi et al., 2018; Wilkins et al., 2019). This so-called symmetry-adapted GPR (SA-GPR) approach was used to paracetamol system leading to a covariant predictions(Raimbault et al., 2019). The same is possible for the class of ML algorithms based on NNs as shown for neuroevolution potential (NEP) scheme(Fan et al., 2021), which was modified into tensorial-NEP approach(Xu et al., 2024) or embedded atom NN (EANN)(Zhang et al., 2019) transformed into tensorial EANN,(Zhang et al., 2020a) outperforming SA-GPR. Other examples introduced new rotationally covariant architectures(Anderson et al., 2019), equivariant message passing NN(Schütt et al., 2021) or tensor field NN(Thomas et al., 2018). The NN-based approaches proved to be excellent tools for predicting infrared and Raman spectra(Chen et al., 2024b; Zhang et al., 2020b), even under the influence of external fields(Zhang and Jiang, 2023).
When considering the kernel ridge regression (KRR) together with a global molecular descriptor, an alternative to these previous approaches is to rotate the molecular geometry to some reference frame(Bereau et al., 2015; Liang et al., 2017; Hu and Huo, 2023). Although such a rotation is straightforward to implement, it cannot be used for flexible or dissociative systems(Grisafi et al., 2018). Here we built upon this rotational idea to propose a three-step procedure to maintain proper covariance. Our procedure is similar in spirit to the one employed by Hu and Huo Hu and Huo (2023), however, we define the reference frame by principal axes of the molecular tensor of inertia rather than by an arbitrarily selected triple of atoms, which should remove a potential source of ambiguity. We demonstrate this new method (named rotate-predict-rotate for reason that will be clear in the next section) on 1,2-dichloroethane, for which new descriptors were introduced improving the overall test set root mean square error (RMSE) values for both dipole moment and polarizability components.